OpenStudy (anonymous):

What is the sum of the angle measures in a polygon that has 20 sides? 2,000 degrees 3,240 degrees 3,600 degrees 4,000 degrees

OpenStudy (anonymous):

Using inductive logic, we can find a pattern: 3 sides: 180 degrees 4 sides: 360 degrees 5 sides: 540 degrees Therefore, the sum of all angles in a polygon is (#sides-2)*180. (20-2)*180=18*180=3240 degrees=option 2